Bromine has two naturally occurring isotopes (Br?79 and Br?81) and has an atomic mass of 79.904 amu. The mass of Br?81 is 80.9163 amu, and its natural abundance is 49.31%.Calculate the mass of Br?79..

Br

Express your answer using four significant figures.

Calculate the natural abundance of Br?79.

Express your answer using four significant figures.
Answer #1

The natural abundance would be 100% - 49.31% = 50.69%


79.904amu = (80.9163)(.4931) + (X)(.5069)

solve for X......mass of Br-79 = 78.922 amu
Answer #2

if Br-81 is 80.9163 amu , and its natural abundance is 49.31%
then
the natural abundance of Br-79 is the rest out of 100 = 50.69%

==============

Calculate the mass
0.5069 (X) + 0.4931(80.9163) = 79.904

0.5069 (X) + 39.8998 = 79.904

0.5069 (X) = 79.904 - 39.8998

0.5069 (X) = 40.0042

X = 78.919